# A Holistic Evaluation of Piano Sound Quality

### v2

@@
Abstract: This paper aims to develop a holistic evaluation method for piano sound quality to assist in purchasing decisions. Unlike previous studies that focused on the effect of piano performance techniques on sound quality, this study evaluates the inherent sound quality of different pianos. To derive quality evaluation systems, the study uses subjective questionnaires based on a piano sound quality dataset. The method selects the optimal piano classification models by comparing the fine-tuning results of different pre-training models of convolutional neural networks (CNN). To improve the interpretability of the models, the study applies equivalent rectangular bandwidth (ERB) analysis. The results reveal that musically trained individuals are better able to distinguish between the sound quality differences of different pianos. The best fine-tuned CNN pre-trained backbone achieves a high accuracy of 98.3% as the piano classifier. However, the dataset is limited, and the audio is sliced to increase its quantity, resulting in a lack of diversity and balance, so we use focal loss to reduce the impact of data imbalance. To optimize the method, the dataset will be expanded, or few-shot learning techniques will be employed in future research.
@@
The sound quality of a piano is of paramount importance for beginners who are looking to purchase one. However, assessing sound quality accurately can be difficult for those without musical training. To aid novice piano buyers, this paper presents a method for evaluating piano sound quality using a convolutional neural networks (CNN) classifier that is incorporated into a mobile application. The proposed method analyzes sound recordings of struck piano keys and generates a sound-quality score, which can help users make more informed purchasing decisions. Moreover, the method can estimate piano prices based on sound quality scores, thereby offering users more reasonable recommendations. Thus, this study has practical implications for improving the piano purchasing process for novices.
@@
To conduct both subjective and objective experiments, a substantial amount of data is required. This part presents the dataset utilized in the experiments, along with a description of the data pre-processing procedures. The dataset and pre-processing methods were essential for conducting the experiments.
@@
The sound sources used in this study were obtained from the CCMUSIC [^3] https://huggingface.co/ccmusic-database platform, a music data-sharing platform in China designed for musicology research. The experiment used a subset of the multifunctional music database, known as the “Piano Sound Quality Dataset”, whose first seven piano classes consist of 12 audio files (in .wav/.mp3/.m4a format) representing seven types of pianos located in the Chinese Conservatory of Music’s piano rooms. These pianos are the Pearl River upright piano (PearlRiver), YOUNG CHANG upright piano (YoungChang), Steinway theater piano (Steinway-T), HSINGHAI upright piano (Hsinghai), KAWAI upright piano (Kawai), Steinway piano (Steinway), and KAWAI grand piano (Kawai-G). The dataset also contains 1,320 single-tone audio files (in .wav/.mp3/.m4a format), resulting in a total of 1,332 files. Moreover, the dataset includes a “Subjective Questionnaire for Evaluating Piano Sound Quality” (in .xls format) that presents scoring data from 29 individuals who participated in the subjective evaluation of piano sound quality. However, the authors did not use the original questionnaire form and results in the dataset but instead conducted their questionnaire. Table 1 provides details of the dataset used.
@@
Steinway-T Steinway-Theater 88-note WAV file 11:29
@@
Kawai-G KAWAI-Grand 88-note WAV file 10:00
@@
Piano Name PearlRiver YoungChang Steinway-T Hsinghai Kawai Steinway Kawai-G

### v3

@@
Abstract: This paper aims to develop a holistic evaluation method for piano sound quality to assist in purchasing decisions. Unlike previous studies that focused on the effect of piano performance techniques on sound quality, this study evaluates the inherent sound quality of different pianos. To derive quality evaluation systems, the study uses subjective questionnaires based on a piano sound quality dataset. The method selects the optimal piano classification models by comparing the fine-tuning results of different pre-training models of convolutional neural network (CNN). To improve the interpretability of the models, the study applies equivalent rectangular bandwidth (ERB) analysis. The results reveal that musically trained individuals are better able to distinguish between the sound quality differences of different pianos. The best fine-tuned CNN pre-trained backbone achieves a high accuracy of 98.3% as the piano classifier. However, the dataset is limited, and the audio is sliced to increase its quantity, resulting in a lack of diversity and balance, so we use focal loss to reduce the impact of data imbalance. To optimize the method, the dataset will be expanded, or few-shot learning techniques will be employed in future research.
@@
The sound quality of a piano is of paramount importance for beginners who are looking to purchase one. However, assessing sound quality accurately can be difficult for those without musical training. To aid novice piano buyers, this paper presents a method for evaluating piano sound quality using a convolutional neural network (CNN) classifier that is incorporated into a mobile application. The proposed method analyzes sound recordings of struck piano keys and generates a sound-quality score, which can help users make more informed purchasing decisions. Moreover, the method can estimate piano prices based on sound quality scores, thereby offering users more reasonable recommendations. Thus, this study has practical implications for improving the piano purchasing process for novices.
@@
To conduct the questionnaire survey and comparative experiments, a substantial amount of data is required. This part presents the dataset utilized in the experiments, along with a description of the data pre-processing procedures. The dataset and pre-processing methods were essential for conducting the experiments.
@@
The sound sources used in this study were obtained from the CCMUSIC [^3] https://huggingface.co/ccmusic-database platform, a music data-sharing platform in China designed for musicology research. The experiment used a subset of the multifunctional music database, known as the “Piano Sound Quality Dataset”, whose first seven piano classes consist of 12 audio files (in .wav/.mp3/.m4a format) representing seven types of pianos located in the Chinese Conservatory of Music’s piano rooms. These pianos are the Pearl River upright piano (PearlRiver), YOUNG CHANG upright piano (YoungChang), Steinway theater piano (Steinway-T), HSINGHAI upright piano (Hsinghai), KAWAI upright piano (Kawai), Steinway piano (Steinway), and KAWAI grand piano (Kawai-G). The dataset also contains 1,320 single-tone audio files (in .wav/.mp3/.m4a format), resulting in a total of 1,332 files. Moreover, the dataset includes a “Subjective Questionnaire for Evaluating Piano Sound Quality” (in .xls format) that presents scoring data from 29 individuals who participated in the subjective evaluation of piano sound quality. However, we did not use the original questionnaire form and results in the dataset, but instead conducted our own questionnaire. Table 1 provides details of the dataset in use.
@@
Steinway-Theater Steinway-T 88-note WAV file 11:29
@@
KAWAI-Grand Kawai-G 88-note WAV file 10:00
@@
Piano label PearlRiver YoungChang Steinway-T Hsinghai Kawai Steinway Kawai-G